OpenStudy (anonymous):

Consider the equation 4x + 6y = 48. Part 1: On your own paper, graph this equation using the slope-intercept method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. Part 2: On your own paper, graph this equation using the intercepts method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar.

hero (hero):

Hints: Part 1: Isolate y: y = 8 - (2x)/3 8 is the y-intercept, which means the first point is @ (0,8) since x = 0. This is your starting point. Next set x = 1 and and solve for y to get the second point. So far you have (1,y). Next set x =2 and solve for y to get the third point. So far you have (2,y). Once you have three points, draw a line through them. Don't forget to label the line. Part 2: When solving using the intercepts method, set x = 0, then solve for y. The setup looks like this: 4(0) + 6y = 48 Solving for y gives you the y coordinate of the graph. Since x = 0, so far you have (0,y). Find y to get the y coordinate. Next set y = 0 to get the other point. The setup looks like this: 4x + 6(0) = 48 Solving for x give you the x coordinate of the graph. So far you have (x,0). Once you plot those two points, you can draw a line through them and you'll have the graph of the line.

OpenStudy (anonymous):

thank you